# Emergent Dynamics in Neural Cellular Automata

### v2

@@ Introduction
Cellular Automata’s (CA) ability to create complex patterns has fascinated researchers for many years (Von Neumann et al.,, 1966; Gardner,, 1970; Wolfram and Gad-el Hak,, 2003). Neural Cellular Automata (NCA [^1] In this paper, we use NCA to refer to both Neural Cellular Automata and Neural Cellular Automaton.), a more recent development (Mordvintsev et al.,, 2020), introduces a trainable variant of conventional CA. The emergence of complex and life-like behaviors, such as spontaneous motion and self-regeneration from simple local update rules, makes NCA a compelling candidate for studying artificial life. NCAs have been applied to a variety of problems, including modeling morphogenesis (Mordvintsev et al.,, 2020; Sudhakaran et al.,, 2021, 2022), texture synthesis (Niklasson et al., 2021b, ; Pajouheshgar et al.,, 2023, 2024), and collective modeling of intelligent behavior (Randazzo et al.,, 2020; Nadizar et al.,, 2022), demonstrating the effectiveness of NCAs in modeling complex biological and natural processes.

### v3

@@ Introduction
Cellular Automata’s (CA) ability to create complex patterns has fascinated researchers for many years (Von Neumann et al.,, 1966; Gardner,, 1970; Wolfram and Gad-el Hak,, 2003). Neural Cellular Automata (NCA [^1] We use NCA to refer to both Neural Cellular Automata and Neural Cellular Automaton.), a more recent development (Mordvintsev et al.,, 2020), introduces a trainable variant of conventional CA. The emergence of complex and life-like behaviors, such as spontaneous motion and self-regeneration from simple local update rules, makes NCA a compelling candidate for studying artificial life. NCAs have been applied to a variety of problems, including modeling morphogenesis (Mordvintsev et al.,, 2020; Sudhakaran et al.,, 2021, 2022), texture synthesis (Niklasson et al., 2021b, ; Pajouheshgar et al.,, 2023, 2024), and collective modeling of intelligent behavior (Randazzo et al.,, 2020; Nadizar et al.,, 2022), demonstrating the effectiveness of NCAs in modeling complex biological and natural processes.